OpenStudy (anonymous):

Solve the equation Check the solution. −2/x+4 = 4/x+3

OpenStudy (johnweldon1993):

\[\large \frac{-2}{x + 4} = \frac{4}{x + 3}\] like that right?

OpenStudy (anonymous):

\[\frac{ -2 }{ x+4 } = \frac{ 4 }{ x+3 }\]

OpenStudy (anonymous):

yes!

OpenStudy (johnweldon1993):

Alright lol....well we can begin by cross multiplying!

OpenStudy (anonymous):

everyone alsways wants to solve the cute girls question lol

OpenStudy (anonymous):

sooo multiply -2 and what? lol

OpenStudy (johnweldon1993):

 like that

OpenStudy (johnweldon1993):

So it would be -2 times the (x + 3) and also the 4 times the (x + 4) so now we will have \[\large -2(x + 3) = 4(x + 4)\] right?

OpenStudy (anonymous):

ohhh. okay okay i see. one sec

OpenStudy (anonymous):

okay whats next

OpenStudy (johnweldon1993):

Well now we solve for 'x' Start by distributing the -2 and the 4 into those parenthesis...

OpenStudy (anonymous):

ummm....x+-6 = x+24

OpenStudy (anonymous):

??

OpenStudy (johnweldon1993):

Not quite....so we had \[\large -2(x + 3) = 4(x + 4)\] right? so what we do...is distribute the -2 into those parenthesis... We take the -2....and multiply it by every term inside the parenthesis...so we will have \[\large -2x - 6 = 4(x + 4)\] make sense? and now we do the same thing for the 4...

OpenStudy (anonymous):

ohh duh. lol so -2x-6 = 4x+24

OpenStudy (johnweldon1993):

Hmm....4 times 4 is what again?

OpenStudy (anonymous):

omg..wow. 16

OpenStudy (johnweldon1993):

haha you're doing good ;P Alright great! so we have \[\large -2x - 6 = 4x + 16\] Now to combine like terms...we want all the terms with an 'x' on 1 side of the equation....and the other ones on the other side

OpenStudy (johnweldon1993):

We can do that by adding 2x to both sides of the equation....and then subtracting 16 from both sides of the equation... what do we have after doing that?

OpenStudy (anonymous):

divide -22by6

OpenStudy (johnweldon1993):

Yes! which is?

OpenStudy (anonymous):

-3.7

OpenStudy (anonymous):

i need it to be an improper fraction

OpenStudy (johnweldon1993):

Great!...so we know that \[\large x = -3.7\] now we check it... And oh okay...then -22/6 it is....but that can be simplified a bit right?

OpenStudy (anonymous):

wait.. how did you get the improper fraction?

OpenStudy (johnweldon1993):

Remember how you said the next step would be "divide -22 by 6" ? :)

OpenStudy (johnweldon1993):

That is the fraction... \[\large x = \frac{-22}{6}\]

OpenStudy (anonymous):

yes i know but how to you turn -3.7 into an improper fraction...

OpenStudy (anonymous):

i understand that -22/6 IS the improper fraction. but how did you get that

OpenStudy (johnweldon1993):

I'm not sure what you mean... alright...lets put it like this... We knew to solve for 'x'....we would divide -22 by 6 When you DO divide -22 by 6...you get the -3.7 \(\large \frac{-22}{6} = \)improper fraction version when you actually DO -22 divided by 6...you get \(\large -3.7 =\) decimal version

OpenStudy (johnweldon1993):

the only thing that we did...was go back 1 step... instead of actually doing the calculation of -22/6 (which comes out to the -3.7 you said) we just left it as the -22/6

OpenStudy (anonymous):

ohhhh >_< okay okay

OpenStudy (johnweldon1993):

lol don't worry about it hun you're fine :) alright...alright....we have \[\large x = \frac{-22}{6}\] right? but...that can be simplified a bit (not calculated...but simplified)

OpenStudy (anonymous):

-11/3

OpenStudy (anonymous):

and thats my answer. right?

OpenStudy (johnweldon1993):

Perfect! that is your improper fraction...and that is your answer :)

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

im gonna tag you in the question